Prince Corporation acquired 100 percent of Sword Company on January 1, 20X7, for $195,000. The trial balances for the two companies on December 31, 20X7, included the following amounts:

Prince Corporation Sword Company
Item Debit Credit Debit Credit
Cash $ 83,000 $ 31,000
Accounts Receivable 67,000 72,000
Inventory 177,000 104,000
Land 81,000 26,000
Buildings and Equipment 491,000 159,000
Investment in Sword Company 255,000
Cost of Goods Sold 491,000 253,000
Depreciation Expense 21,000 11,000
Other Expenses 66,000 66,000
Dividends Declared 52,000 26,000
Accumulated Depreciation $ 143,000 $ 55,000
Accounts Payable 64,000 30,000
Mortgages Payable 185,000 108,000
Common Stock 287,000 45,000
Retained Earnings 324,000 91,000
Sales 695,000 419,000
Income from Sword Company 86,000
$ 1,784,000 $ 1,784,000 $ 748,000 $ 748,000


Additional Information

  1. On January 1, 20X7, Sword reported net assets with a book value of $136,000. A total of $26,000 of the acquisition price is applied to goodwill, which was not impaired in 20X7.
  2. Sword’s depreciable assets had an estimated economic life of 11 years on the date of combination. The difference between fair value and book value of tangible assets is related entirely to buildings and equipment.
  3. Prince used the equity-method in accounting for its investment in Sword.
  4. Detailed analysis of receivables and payables showed that Sword owed Prince $25,000 on December 31, 20X7.


Required:
a. Prepare all journal entries recorded by Prince with regard to its investment in Sword during 20X7. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)



b. Prepare all consolidating entries needed to prepare a full set of consolidated financial statements for 20X7. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)



c. Prepare a three-part consolidation worksheet as of December 31, 20X7. (Values in the first two columns (the "parent" and "subsidiary" balances) that are to be deducted should be indicated with a minus sign, while all values in the "Consolidation Entries" columns should be entered as positive values. For accounts where multiple adjusting entries are required, combine all debit entries into one amount and enter this amount in the debit column of the worksheet. Similarly, combine all credit entries into one amount and enter this amount in the credit column of the worksheet.)
Answer #1

Part A

No.

Event

General journal

Debit

Credit

A

1

Investment in Sword Company

195000

Cash

195000

B

2

Investment in Sword Company (419000-253000-11000-66000)

89000

Income from Sword Company

89000

C

3

Cash

26000

Investment in Sword Company

26000

D

4

Income from Sword Company (195000-136000-26000)/11

3000

Investment in Sword Company

3000

Part B

No.

Event

General journal

Debit

Credit

A

1

Common stock

45000

Retained earnings

91000

Income from Sword Company

89000

Dividends declared

26000

Investment in Sword Company (balancing figure)

199000

(To record basic consolidation entry)

B

2

Depreciation expense

3000

Income from Sword Company

3000

(To record amortized excess value reclassification entry)

C

3

Buildings and equipment

33000

Goodwill

26000

Accumulated depreciation

3000

Investment in Sword Company (balancing figure)

56000

(To record the excess value (differential) reclassification entry)

D

4

Accounts payable

25000

Accounts receivable

25000

(To record entry to eliminate the intercompany accounts)

E

5

Accumulated depreciation (55000-(4*3000))

43000

Buildings and equipment

43000

(To record the optional accumulated depreciation consolidation entry)

Fair value = 195000

Book value = 136000

Excess value = 59000

Assigned to goodwill = 26000

Excess assigned to building and equipment = 59000-26000 = 33000

Amortization of excess assigned to building and equipment = 33000/11 = $3000

PRINCE CORPORATION AND SUBSIDIARY

Consolidated Financial Statements Worksheet

December 31, 20X7

Consolidation Entries

Prince Corp.

Sword Co.

Dr.

Cr.

Consolidated

Income statement

Sales

695000

419000

1114000

Less: COGS

(491000)

(253000)

-744000

Less: depreciation expense

(21000)

(11000)

3000

-35000

Less: other expenses

(66000)

(66000)

-132000

Income from Sword Co.

86000

89000

3000

0

Net income

203000

89000

92000

3000

203000

Statement of retained earnings

Beginning balance

324000

91000

91000

324000

Net income

203000

89000

92000

3000

203000

Less: dividends declared

(52000)

(26000)

26000

(52000)

Ending balance

475000

154000

183000

29000

475000

Balance sheet

Assets

Cash

83000

31000

114000

Accounts receivable

67000

72000

25000

114000

Inventory

177000

104000

281000

Land

81000

26000

107000

Buildings & equipment

491000

159000

33000

43000

640000

Less: Accumulated depreciation

-143000

-55000

43000

3000

-158000

Investment in Sword Co.

255000

255000

0

Goodwill

26000

26000

Total Assets

1011000

337000

102000

326000

1124000

Liabilities & Equity

Accounts payable

64000

30000

25000

69000

Mortgages payable

185000

108000

293000

Common stock

287000

45000

45000

287000

Retained earnings

475000

154000

183000

29000

475000

Total Liabilities & Equity

1011000

337000

253000

29000

1124000

Investment in Sword Co. = 195000+89000-26000-3000 = 255000